OpenStudy (anonymous):

solve for x and y, using substitution. 2x-y=8 x+2y=9

OpenStudy (anonymous):

Pick one of the two equations and solve for either x or y. For example: x + 2y = 9 x = 9 - 2y Now plug 9 - 2y into the other equation in place of x. You can then solve that equation for y and, once you've done that, use the value of y to calculate x.

OpenStudy (anonymous):

solve for y in the first equation. so y=2x-8. then enter it into the second equation. so x+2(2x-8)=9. solve for x and then enter your x value into the equation x+2y=9 to find your y value.

OpenStudy (anonymous):

x+4x-16=9 ?

OpenStudy (anonymous):

Would x=5

OpenStudy (anonymous):

yup. now enter that into x+2y=9 to find your y value.

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

yup thats it

OpenStudy (anonymous):

can you help me with some more

OpenStudy (anonymous):

ya sure give me another

OpenStudy (anonymous):

y=2x+1 and y=3x-7, I think i found out x

OpenStudy (anonymous):

ok so do 3x-7=2x+1 x will come out to x=8

OpenStudy (anonymous):

now enter that value into one of the equations given to find y

OpenStudy (anonymous):

1x-7=1?

OpenStudy (anonymous):

yes. add the 7 to the other side so that x is by itself and you get x=8

OpenStudy (anonymous):

Nevermind its 17, right?

OpenStudy (anonymous):

is it 8?

OpenStudy (anonymous):

ur killin me dude. 7+1=8. haha maybe a typo or somthin

OpenStudy (anonymous):

ya haha

OpenStudy (anonymous):

y = 17 my bad. nice

OpenStudy (anonymous):

alright thanks, I have like 3 more problems. sorry im really bad at math

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x+y=4 and -6x-2y=12

OpenStudy (anonymous):

the easiest way would be to solve for the y in the first equation. so y=4-3x

OpenStudy (anonymous):

now enter that into the second equation

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-6x-2(4-3x)=12 solve for x

OpenStudy (anonymous):

x-8=12 ?

OpenStudy (anonymous):

well since -6x-8+6x=12. the x cancels out so you are left with -8=12. this means that there is no solution to the problem

OpenStudy (anonymous):

the answer is "no solution"

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

you got that.

OpenStudy (anonymous):

did u have one more

OpenStudy (anonymous):

2 more, y=x+1 and x=y-1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since u already have an x or y solved for, pick one to put into the other equation. so im gonna enter the y one in for the y in the x equation. so x=(x+1)-1.

OpenStudy (anonymous):

2= -x ?

OpenStudy (anonymous):

huh well sorry but for some reason this one is stumping me. give me the next one and i will come back to it

OpenStudy (anonymous):

ok, 8x-6y = 22 and x-3y=-4

OpenStudy (anonymous):

ok so solve for x in the second one. x=3y-4. enter it into the first one so 8(3y-4)-6y=22. solve for y.

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

yup nice. now enter that back in to get x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x= 9-4 ?

OpenStudy (anonymous):

yup exactly

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

yup now for that last one. i figured it out by just looking at it. x=-2 and y =-1. i cant really explain it though. when i try to solve for it i must be making a lil mistake. but if u enter those nunbers in they work so sorry i cant walk ya through it

OpenStudy (anonymous):

its okay and thanks

OpenStudy (anonymous):

ok have a good night

OpenStudy (anonymous):

you too